ABC Ltd reported that for the financial year ended June 30, 2020, depreciation expense was $78,000 and COGS was $847,000. They also reported a Loss on Sale of $25,000 on equipment which had a book value of $60,000.

Additionally their Balance Sheet at June 30, 2019 reported Inventory at $54,000, Plant and Equipment (net) of $627,000 and Accounts Payable of $29,000. The figures for June 30, 2020 were Inventory, $61,000, Plant and Equipment (net) at $710,000 and Accounts Payable $32,000.

  1. How much cash was paid for inventory purchases during the year?
  2. What was the cost of the new plant acquired?

(State any assumptions you make)
Answer #1

a) Cash paid for inventory purchase = 847000+7000-3000 = 851000

b) Purchase of plant and equipment = Ending value+Book value of sale of equipment-Beginning value

= 710000+60000-627000

Purchase of plant and equipment = 143000

Cost of new plant acquired = $143000